OpenStudy (anonymous):

write a function rule that represents each sentence for 12 more than the quotient of a number T and 7 is v

OpenStudy (anonymous):

\[number /divisor =quotient +remainder ..........T/7 =v+12\]

OpenStudy (anonymous):

unless the and means plus then T+7=v+12